OpenStudy (anonymous):

What is the solution to the rational equation x over 2x plus 1 minus 1 over 4 equals 2 over 2x plus 1 ? x = 9 over 2 x = 7 over 2 x = 2 x = 3 i will reward a medal

OpenStudy (anonymous):

\[\frac{ x }{ 2x+ 1 }-\frac{ 1 }{ 4 }=\frac{ 2 }{ 2x +1 }\] is the equation and the answer options are \[x=\frac{ 9 }{ 2 }\]\[x=\frac{ 7 }{ 2 }\] x=2 x=3